The cost of a chair and a table is Rs.2000 .The cost of 5 chairs and some tables is 16200, if number of table and chair are interchanged and 5% discount is allowed .I have to pay Rs.11210.Find number of tables I initially purchase? Find the total cost of seven chairs and six tables?
Answer by stanbon(75887)

The cost of a chair and a table is Rs.2000 .The cost of 5 chairs and some tables is 16200, if number of table and chair are interchanged and 5% discount is allowed .I have to pay Rs.11210.Find number of tables I initially purchase? Find the total cost of seven chairs and six tables?
-----------------
c + t = 2000
5c + xt = 16200
xc + 5t = 11210
-------------------
(5-x)c + (x-5)t = 4900
----
Substitute for c (or for t, and solve for the other)
Then you can figure out the cost of 7c + 6t.
----
Cheers,
Stan H.